News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Sunday 2 April 2000 3.00pm EDT

During the last week, Chandra completed the observing schedule as planned.

The observatory passed through the final (13th) eclipse of the Spring 2000 season with nominal power and thermal performance. The eclipse was a short duration eclipse (40 min) and occurred at 2:43am on 3/28, lasting for 32.7 minutes.

A real-time procedure was completed on 3/29 to uplink a software constant patch to the Electrical Power System to configure for the solstice period (March-September).

The scheduled observations for the next week are:

------------------------------------------
MITG J0414+0534        ACIS-S          Apr  2

B2 0755+37             ACIS-S

W3B                    ACIS-I

BAL 1044+3656          ACIS-S          Apr  3

Q1331+170              ACIS-S

MKW 3S                 ACIS-I

1735-269               ACIS-S          Apr  4

E0102-72.3             ACIS-I

1E 0056.8-7154         ACIS-S

Radiation Belt

MCG-6-30-15            ACIS-S/HETG     Apr  5

RX J1347-114           ACIS-S          Apr  6

UM 425                 ACIS-S

NGC 4406               ACIS-S

Radiation Belts                        Apr  7

Zeta Ori               ACIS-S/HETG

4U 1538-52             ACIS-I          Apr  8

RE 0751+14             HRC-S/LETG      Apr  9

Zeta Ori               ACIS-S/HETG

1E 0056.8-7154         ACIS-S

Radiation Belts

------------------------------------------

Note that a major radio flare was detected from the the X-ray source Cygnus X-3 on April 1. The source has been accepted as a Target Of Opportunity (TOO) and rescheduling work began on 4/2. The next status report will include schedule revisions made to to accommodate the Cygnus X-3 observations.

The spacecraft subsystems continued to operate nominally.

Return to Status Reports